Is pair of lines parallel, perpendicular, or neither?
f(x)=-4x+7
g(x)=7x+3

Found 2 solutions by solver91311, richwmiller:
Answer by solver91311(24713)



Both are in slope-intercept form so that you can determine the slope of each by inspection. If the slopes are equal, the graphs are parallel (or possibly coincident). If the slopes are negative reciprocals, then the graphs are perpendicular. Otherwise, neither.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by richwmiller(17219)

Not parallel nor perpendicular. They intersect at x=4/11 and y=61/11